OpenStudy (anonymous):

The first term of an arithmetic sequence is -5, and the tenth term is 13. Find the common difference. 8/9 1.8 2

OpenStudy (mathstudent55):

Let's look at what the terms of an arithmetic sequence with first term \(a_1\) and common difference \(d\) look like: \(a_1 = a_1\) \(a_2 = a_1 + d\) \(a_3 = a_1 + 2d\) \(a _4 = a_1 + 3d\) \(...\) \(a_n = a_1 + (n - 1)d\) You know \(a_n\), \(a_1\), and \(n\). Solve for \(d\).